OpenStudy (anonymous):

\[\frac{ \sin ^{2}x-\sin x-2 }{ \sin x-2 }? \] what is equivalent

OpenStudy (mathstudent55):

The numerator is a quadratic equation in \(\sin x\). Try factoring the numerator.

OpenStudy (mathstudent55):

If factoring the numerator directly looks confusing, then use a substitution. Let \(u = \sin x\) Then the numerator is \(u^2 - u - 2\) Can you factor that quadratic equation in u?

OpenStudy (jdoe0001):

\(\bf \cfrac{ \sin ^{2}(x)-\sin (x)-2 }{ \sin (x)-2 }\implies \cfrac{ [{\color{brown}{ sin(x)}}]^2-{\color{brown}{ sin (x)}}-2 }{ {\color{brown}{ sin (x)}}-2 }\implies \cfrac{ [{\color{brown}{ u}}]^2-{\color{brown}{ u}}-2 }{ {\color{brown}{ u}}-2 }\) notice what mathstudent55 suggested

OpenStudy (anonymous):

sin^2x?

OpenStudy (anonymous):

sin x^2 - 1?

OpenStudy (mathstudent55):

No. Can you factor \(u^2 - u - 2 \) ?

OpenStudy (anonymous):

no?

OpenStudy (mathstudent55):

If you are asking if that is a final answer, the answer is no.

OpenStudy (anonymous):

ok so what is the answer?

OpenStudy (mathstudent55):

I can give you the answer, but that only shows that I know ho to solve it, not that you learned anything. I am trying to help you by breaking down this problem into simple steps, but you haven't answered any of the questions I asked.

OpenStudy (mathstudent55):

Factor: \(u^2 - u - 2\)

OpenStudy (anonymous):

(u-2)(u+1)

OpenStudy (anonymous):

now what?

OpenStudy (mathstudent55):

Great Now since we substituted u for sin x, let's substitute sin x back in for u.

OpenStudy (anonymous):

sin x -2 and sin x +1

OpenStudy (mathstudent55):

This is everything we have done so far: \(\dfrac{\sin^2 x - \sin x - 2}{\sin x - 2} \) Let \(u = \sin x\) \(=\dfrac{u^2 - u - 2}{\sin x - 2} \) \(= \dfrac{(u - 2)(u + 1)}{\sin x - 2} \) Now we substitute \(\sin x\) back in for u: \(=\dfrac{(\sin x - 2)(\sin x + 1)}{\sin x - 2} \)

OpenStudy (mathstudent55):

Now what cancels out?

OpenStudy (anonymous):

sin x - 2

OpenStudy (jdoe0001):

yeap thus \(\bf \cfrac{ \sin ^{2}(x)-\sin (x)-2 }{ \sin (x)-2 }\implies \cfrac{ [{\color{brown}{ sin(x)}}]^2-{\color{brown}{ sin (x)}}-2 }{ {\color{brown}{ sin (x)}}-2 }\implies \cfrac{ [{\color{brown}{ u}}]^2-{\color{brown}{ u}}-2 }{ {\color{brown}{ u}}-2 } \\ \quad \\ \cfrac{\cancel{ ({\color{brown}{ u}}-2)}({\color{brown}{ u}}+1)}{\cancel{ {\color{brown}{ u}}-2}}\implies \cfrac{{\color{brown}{ sin(x)}}+1}{1}\implies sin(x)+1\)

OpenStudy (anonymous):

so sin x + 1 is the answer

OpenStudy (mathstudent55):

Correct

OpenStudy (mathstudent55):

\(=\dfrac{(\cancel{\sin x - 2})(\sin x + 1)}{\cancel{\sin x - 2}~~1} \) \(=\sin x + 1\)